OpenStudy (anonymous):

(7 + 2x) 2x = (23 + x) x. solve for x

OpenStudy (shaik0124):

14x+4x2=23x+x2 4x2-x2=23x-14x 3x2=9x 3x=9 x=9/3 x=3

OpenStudy (anonymous):

Thank you so much!!!